OpenStudy (anonymous):

Does this sum converge? n=1 to infinity of (1/3^n) - (5/sqrt(n))

OpenStudy (sirm3d):

\[\sum_{1}^{\infty} \frac{ 1 }{ 3^n }-\frac{ 5 }{ \sqrt{n} }\] is this the problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sum (1/n) diverges i/sqrt(n) = sqrt(n)/n > 1/n so sum(1/sqrt(n)) diverges, and the whole expression will.